OpenStudy (bluebeta):

The angle of elevation to the top of a skyscraper is measured to be 2° from a point on the ground 1 mile from the building. How tall is the skyscraper? My work is in the comments.

OpenStudy (bluebeta):

I got 5276.78 ft. tall. I changed 1 mile=5280 ft Cos2=x/5280 Cos2*5280 Cos2=.9993908 .9993908*5280=5276.78 ft @colinvda I seen you look at this.

OpenStudy (anonymous):

Hey, no stalking me ;) You have the right idea, but think about the x/y/r fractions I gave you in your other question. Are you sure you should be using cos?

OpenStudy (anonymous):

Try drawing the question. Label the 3 sides of the triangle and the 2°

OpenStudy (bluebeta):

@colinvda ok ill try that

OpenStudy (radar):

The hypotenuse is the longest side of a right triangle. From this we know the distance is in excess of one mile. A picture is worth the effort.

OpenStudy (bluebeta):

@radar @colinvda okay.. i don't get how to fix it though..